Circular No. 7211

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990627 J. M. Muller, Space Research Organization of the Netherlands (SRON),

Utrecht, and BeppoSAX, Rome; E. Costa and G. Gandolfi, Istituto Astrofisica Spaziale, CNR, Rome; M. Dadina, A. Paolino, R. Ricci, and A. Tesseri, BeppoSAX, Rome; and J. Heise and J. J. M. in ’t Zand, SRON, report on behalf of the BeppoSAX team: “A γ-ray burst with x-ray counterpart has been detected by the BeppoSAX Gamma Ray Burst Monitor (GRBM) and Wide Field Camera 2 (WFC2) on June 27.20894 UT. A preliminary analysis of the GRBM data reveals a multiple-peak γ-ray burst lasting ∼ 50 s with a peak intensity of 280 counts/s in 1 s timebins in the energy band 40–700 keV. A simultaneous x-ray counterpart was observed by the WFC2 with a duration of ∼ 60 s, a complex lightcurve, and peak and mean fluxes of ∼ 0.3 and 0.1 Crab (2-9 keV), respectively. The position of the counterpart resulting from an off-line analysis is α = 1^{h}48^{m}25^{s}, δ = −77^{o}05′.3 (equinox 2000.0), with an error radius of 3′. A follow-up observation with the BeppoSAX narrow-field instruments is in progress since June 27.55. Observations in all bands are solicited. The SIMBAD database

(http://cdsweb.u-strasbg.fr/Simbad.html#inter) mentions a star of

mag 9.9 within the error circle.”

SUPERNOVA 1999cw IN MCG –01-02-001 R. Johnson and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on June 28.5 and 29.5 UT, both with mag ∼ 14.2. The new object is located at α = 0^{h}20^{m}01^{s}.46, δ = −6^{o}20′03′′.6 (equinox 2000.0), which is 21′′.1 east and 1′′.5 south of the nucleus of MCG –01-02-001. A KAIT image of the same field on 1998 Aug. 15.5 (limiting magnitude ∼ 19.0) showed nothing at the position of SN 1999cw.

NOVA IN NGC 6822 G. Bakos, Space Telescope Science Institute, and the PLANET col-

laboration report that 420- and 360-s V -band images of the apparent nova in NGC 6822 (IAUC 7208 and 7209) were taken with the 1.0-m telescope at Sutherland Observatory on June 26 and 28, respectively, yielding the following derived magnitudes (based on USNO-A1.0 R magnitudes): June 26.082 UT, V = 19.0 ± 0.1; 28.074, 19.6 ± 0.1.

1999 June 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green